OpenStudy (anonymous):

-14x +15y = 15 21x -20y = -10

OpenStudy (smokeysthename):

solve by?... elimination? substitution? which one?

OpenStudy (anonymous):

elimination.

OpenStudy (smokeysthename):

okay so you have to multiply the x or y in both equations so they are the same number

OpenStudy (anonymous):

I know :( but I don't nkw what number to multiply.

OpenStudy (smokeysthename):

multiply the top equation by 3 and the bottom equation by 2 and lemme know what you get.

OpenStudy (smokeysthename):

@Rubie you there?

OpenStudy (anonymous):

@SmokeysTheName yupp

OpenStudy (smokeysthename):

okay so what did you get ?

OpenStudy (smokeysthename):

what did you get when you multiplied the top equation by 3 and the bottom equation by 2?

OpenStudy (anonymous):

wait a minute lemme check.

OpenStudy (smokeysthename):

k

OpenStudy (anonymous):

-42x + 45y = 45

OpenStudy (anonymous):

42x - 40y = -20

OpenStudy (smokeysthename):

correct (: now you can combine both equations so it would be -42x+42x+45y-40y=45-20 simplify that and tell me what you get

OpenStudy (anonymous):

5y = 23

OpenStudy (smokeysthename):

no not quite. you're close, on the right side of the equal sign its 45-20 which equals 25, not 23

OpenStudy (smokeysthename):

therefore you have 5y=25 and from there you can solve for y. y=..?

OpenStudy (anonymous):

y = 5

OpenStudy (smokeysthename):

correct (: now plug y back into the original equation that you started with. you can choose either one, you'll get the same answer.

OpenStudy (anonymous):

-16x = -60?

OpenStudy (smokeysthename):

do you chose the -14x+15y=15 equation?

OpenStudy (anonymous):

yes.

OpenStudy (smokeysthename):

how did you get -16x

OpenStudy (anonymous):

well, -14x + 75 = 15

OpenStudy (anonymous):

and then -16x = -60

OpenStudy (smokeysthename):

you dont change -14x . im not sure why you keep changing it. -14x+75=15 subtract 75 from both sides -14x=-60 understand?

OpenStudy (anonymous):

ohhh okokok

OpenStudy (anonymous):

ans is 30/7?

OpenStudy (smokeysthename):

correct (: or if you put it into decimal its x=4.3

OpenStudy (smokeysthename):

good job @Rubie !! (:

OpenStudy (anonymous):

thank oyuu

OpenStudy (smokeysthename):

no problem (: